# Symbiosis: Multi-Adapter Inference and Fine-Tuning

### v2

@@ 4.4. Mixed Fine-tuning and Inference
Configuration Latency
Small & Small 0.30
Small & Large 3.74
Large & Large 6.94
Table 4. vLLM prefill response time for batched inference of small and large requests for Llama2-7b model. Small: sequence length=1, Large: sequence length=512.
@@ 4.5. Opportunistic Batching
In comparison, Table 5 shows the benefit of opportunistic batching for inference requests with different characteristics. Specifically, we create a variety in clients by using different batch sizes (2 to 256) and different LoRA adapters, LoRA 1 and LoRA 4 (from Table 2). Such a setting is typical for a service provider, who expects requests from a variety of clients. In the no-lockstep approach, each request can progress through the model as quickly as possible without having to batch with other requests. However, with a large number of requests, such approach causes the execution of different layers by different requests to be serialized, which increases response time and the lack of batching also reduces inference throughput. In lockstep policy, the base executor waits for all the requests for every base model layer. Because of the differences in client-side computation, namely, attention and adapter calculation, this policy suffers from longer wait times. In contrast, Symbiosis’s opportunistic batching waits only for a pre-determined duration for batching and then proceeds with the accumulated requests. Importantly, the requests batched at first layer are not required to be batched again for the following layers, which makes different rate of execution for different clients possible. The wait time factor can be configured by the service provider to capture the worse tolerable penalty in terms of latency. In this experiment, the request with 256 batch size needs to wait for 50ms at maximum per-iteration.
Throughput Latency Average
(tokens/s) (s) Batch Size
No Lockstep 94 1.02 1
Lockstep 88 1.6 8
Opportunistic Batching 103 0.77 3.7
Table 5. Average response time of 8 inference jobs (token generation) in remote client configuration. The batch sizes (from 2 to 256) and adapter types vary for different clients.
@@ 4.6. Privacy for Multi-Tenancy
Figure 23 demonstrates the effect of privacy on inference performance. We host a client on a separate host, with the assumption that such host is under tenant control. Whereas the base model can be hosted by a service provider. The tenant adds noise to the activation before transmitting them over the network and the noise effect is deducted from the received output. It can be observed that the effect of noise addition and subtraction is minimal. This is because we calculate the noise effect for each layer in advance. Moreover, the network interference is a primary factor in the degraded performance when compared to the communication between the GPUs on the same host. However, even with the multi-tenant setup the performance remains acceptable. Symbiosis uniquely enables this use case where tenant parameters (of the adapter) and activations are protected, while allowing them to leverage a common base model.
@@ Acknowledgment
Thanks to Ka-Ho Chow for his help with the initial prototype.

### v3

@@ 4.5. Opportunistic Batching
In comparison, Table 5 shows the benefit of opportunistic batching for inference workload from Azure Public Dataset (azu, 2025). This is a sample of the traces from multiple LLM inference services in Azure. The dataset contains a trace of inference requests with context sequence lengths of up to 7K tokens and generated sequence lengths of up to 32 tokens. We use Llama2-13B as the base model deployed in sharded local configuration across 8 GPUs. 32 inference clients share the base model for inference (4 clients per GPU). Each client uses a 64-rank LoRA adapter with q and k as the fine-tuned layers. Together all clients replay a portion of the inference workload trace from the Azure Public Dataset. In the no-lockstep approach, each request can progress through the model as quickly as possible without having to batch with other requests. However, with a large number of requests, such approach causes the execution of different layers by different requests to be serialized. The serialization of requests increases latency and the lack of batching reduces inference throughput. In contrast, Symbiosis’s opportunistic batching waits only for a pre-determined duration for batching and then proceeds with the accumulated requests. This improves throughput from better batching. For instance, the generation throughput with opportunistic batching is 28% higher than that with no-lockstep. Note that the throughput improvement in prefill is not as significant as generation from the already large context lengths, which limit the further throughput improvement from batching. Additional benefit of opportunistic batching is that the requests batched at the first layer are not required to be batched again for the following layers, which makes different rate of execution for different clients possible. This behavior is useful in maintaining lower latency.
@@ 4.6. Privacy for Multi-Tenancy
Figure 23 demonstrates the effect of privacy on inference performance of model Llama2-7B. We host a client on a separate host, with the assumption that such host is under tenant control. Whereas the base model can be hosted by a service provider. The tenant adds noise to the activation before transmitting them over the network and the noise effect is deducted from the received output. It can be observed that the effect of noise addition and subtraction is minimal. This is because we calculate the noise effect for each layer in advance. Moreover, the network interference is a primary factor in the degraded performance when compared to the communication between the GPUs on the same host. However, even with the multi-tenant setup the performance remains acceptable. Symbiosis uniquely enables this use case where tenant parameters (of the adapter) and activations are protected, while allowing them to leverage a common base model.
@@ Acknowledgment
Thanks to Ka-Ho Chow for his contribution to the initial prototype, to Ryan Marcus for his guidance as a shepherd, and to our reviewers for their valuable suggestions.